OpenStudy (anonymous):

Solve 8.1x – 3.2 = 7.6x – 2.4.

sam (.sam.):

x=1.6

sam (.sam.):

8.1x-3.2=7.6x-2.4 Move all terms not containing x to the other side of the equation 0.5x-3.2=-2.4 0.5x=0.8 Divide each term in the equation by 0.5. x=1.6

OpenStudy (anonymous):

x = 1/3 x = -1/3 x = -1 x = 1